### v3

# Generalized electrical Lie algebras

Abstract.
We generalize the electrical Lie algebras originally introduced by Lam and Pylyavskyy in several ways. To each Kac-Moody Lie algebra $\mathfrak{g}$ we associate two types (vertex type and edge type) of the generalized electrical algebras. The electrical Lie algebras of vertex type are always subalgebras of $\mathfrak{g}$ and are flat deformations of the nilpotent Lie subalgebra of $\mathfrak{g}$. In many cases including $sl_{n}$, $so_{n}$, and $sp_{2n}$ we find new (edge) models for our generalized electrical Lie algebras of vertex type. Finding an edge model in general is an interesting an open problem.
This work was partially supported by the Simons Foundation Collaboration Grant for Mathematicians no. 636972 (AB)
MSC2020: 14L40, 20G05, 82B20, 05E10
Key words: Electrical Lie Algebra, Electrical Networks, Serre relations
1. Introduction and main results
The electrical Lie algebras were introduced by T. Lam and P. Pylyavskyy in [8] and further studied by Yi Su in [10]. An infinitely generated electrical Lie algebra of type $A$ also appeared in the context of categorification in representation theory [2]. The aim of this paper is to generalize the notion of electrical Lie algebra into a multiparametric family in any given semisimple or Kac-Moody Lie algebra $\mathfrak{g}$, in such a way that they are flat deformations of the nilpotent part $\mathfrak{n}$ of $\mathfrak{g}$. We also construct various embeddings (vertex and edge models) of the electrical Lie algebras into the corresponding Kac-Moody ones.
We start with the definition of the electrical Lie algebra given below which depends on a set of parameters $a_{i}$. If all $a_{i}$ are equal to the imaginary unit these generators in the case of $sl_{n}$ give an important representation of the Temperley-Lieb algebra for the zero value of the loop fugacity parameter ([1]) and they also give a solution to the Zamolodchikov tetrahedral equation [9], [6].
Definition 1.1.
Let $A=(a_{ij})$ be a (generalized, not necessarily symmetrizable) $I\times I$ Cartan matrix and let $\mathfrak{g}=\mathfrak{g}_{A}=<e_{i},f_{i},i\in I>$ be the corresponding Kac-Moody Lie algebra. For any family ${\bf a}=(a_{i},{i\in I})\in\mathbb{C}^{I}$ let $\mathfrak{g}^{({\bf a})}$ be a Lie subalgebra of $\mathfrak{g}$ generated by
$u_{i}:=e_{i}+a_{i}[e_{i},f_{i}]-a_{i}^{2}f_{i}\ ,i\in I.$ (1.1)
We call $\mathfrak{g}^{({\bf a})}$ a generalized electrical Lie algebra of type $\mathfrak{g}$.
In particular, $sl_{2}^{({\bf a})}$ is generated by a single nilpotent $u=\begin{pmatrix}a&1\\ -a^{2}&-a\end{pmatrix}\in sl_{2}$.
This definition makes sense if the ground field $\mathbb{C}$ is replaced with any algebra containing all $a_{i}$, e.g., with $\mathbb{C}[a_{i},i\in I]$ and it is justified by the following.
Theorem 1.2.
For any Kac-Moody Lie algebra $\mathfrak{g}$ and any ${\bf a}\in\mathbb{C}^{I}$ the generalized electrical Lie algebra $\mathfrak{g}^{({\bf a})}$ of type $\mathfrak{g}$ satisfies
$(ad~{}u_{i})^{1-a_{ij}}(u_{j})=-2\delta_{a_{ij},-1}a_{ji}a_{i}a_{j}u_{i}$ (1.2)
for all distinct $i,j\in I$.
It turns out that these relations are defining. This was conjectured for finite types in [8, Conjecture 5.2], which which was partially confirmed in [10] and [3, 4].
Theorem 1.3.
In the notation of Theorem 1.2 the relations (1.2) provide a presentation of $\mathfrak{g}^{({\bf a})}$. Moreover, $\mathfrak{g}^{({\bf a})}$ is a flat deformation of the nilpotent part ${\mathfrak{n}}=<e_{i},i\in I>$ of $\mathfrak{g}$ in the sense that $\mathfrak{g}^{({\bf a})}\cong\mathfrak{n}$ for any ${\bf a}\in\mathbb{C}^{I}$ as a naturally filtered Lie algebra.
We prove Theorems 1.3 in Section 3.2 (along with their generalization to any coefficient algebra $R$ containing all $a_{i}$) by utilizing a mini-theory of flat deformations of associative and Lie algebras which we present for the reader’s convenience in Appendix.
Note that if we view all $a_{i}$ as formal parameters (after replacing $\mathbb{C}$ with $\mathbb{C}[{\bf a}]=\mathbb{C}[a_{i},i\in I]$), then $\mathbb{C}[{\bf a}]\otimes\mathfrak{g}$ is graded by $\mathbb{Z}^{I}_{\geq 0}$ via $\deg a_{i}=\deg e_{i}=-\deg f_{i}$, $i\in I$. In particular, $u_{i}$ is homogeneous with $\deg u_{i}=\deg a_{i}$ and both $\mathfrak{g}^{({\bf a})}$ and $U(\mathfrak{g}^{({\bf a})})$ are $\mathbb{Z}^{I}_{\geq 0}$-graded.
Now we construct a real form $\mathfrak{g}^{({\bf b})}$ for the electrical Lie algebra $sl_{n}^{({\bf b})}$.
Theorem 1.4.
Let $\mathfrak{g}=sl_{n}$. Then $g_{\bf a}u_{i}g_{\bf a}^{-1}=e_{i}+b_{i-1}f_{i-1}$ for $i=1,\ldots,n-1$ in the notation of (1.1) (with the convention $f_{0}=0$, $a_{0}=0$), where we abbreviated $b_{j}:=-a_{j}a_{j+1}$ and $g_{\bf a}:=e^{a_{n-1}f_{n-1}}\cdots e^{a_{2}f_{2}}e^{a_{1}f_{1}}\in SL_{n}$.
We prove Theorem 1.4 in Section 3.3.
Remark 1.5.
This realization with all $a_{i}=\sqrt{-1}$ (i.e., all $b_{j}=1$) was first discovered by T. Lam and A. Postnikov as it is mentioned in [7].
The realization with all $a_{i}$ imaginary is a “real form” of the electric Lie algebra of type $sl_{n}$. We produce more real forms for $\mathfrak{g}^{(\bf a)}$ below and in Section 2.
More generally, for any ${\bf b}=(b_{1},\ldots,b_{n-2})\in\mathbb{C}^{n-2}$ we denote by $sl_{n}^{(\bf b)}$ the Lie subalgebra of $sl_{n}$ generated by $u_{i}:=e_{i}+b_{i-1}f_{i-1}$, $i=1,\ldots,n-1$ (with the convention $b_{0}=0$, $f_{0}=0$ and establish the following
Theorem 1.6.
The Lie algebra $sl_{n}^{(\bf b)}$ has a presentation
$\bullet$ $[u_{i},u_{j}]=0$ if $|i-j|>1$
$\bullet$ $[u_{i},[u_{i},u_{j}]]=-2b_{\min(i,j)}u_{i}$ if $|i-j|=1$
We prove Theorem 1.6 in Section 3.5 as a particular case of Theorem 2.3.
Clearly, if all $b_{i}\neq 0$, Theorem 1.6 follows from Theorem 1.4. Note however, that for $n=5$, $b_{1}\neq 0$, $b_{2}=0$, $b_{3}\neq 0$, such a tuple $(a_{1},a_{2},a_{3},a_{4})$ does not exist and the assertion of Theorem 1.6 does not follow from Theorem 1.4.
Theorem 1.4 suggests an “edge model” of generalized electrical Lie algebras as follows.
Given a Kac-Moody Lie algebra $\mathfrak{g}$ and a tuple ${\bf b}=\{b_{ij}=b_{ji}|a_{ij}=-1\}$, we denote by $\mathfrak{g}^{({\bf b})}$ the Lie algebra generated by $u_{i}$, $i\in I$ subject to
$(ad~{}u_{i})^{1-a_{ij}}(u_{j})+2\delta_{a_{ij},-1}b_{ij}u_{i}=0$ (1.3)
for all distinct $i,j\in I$ where $b_{ij}=0$ unless $a_{ij}=-1$. We refer to $\mathfrak{g}^{({\bf b})}$ as the generalized electrical Lie algebra of the edge type $\mathfrak{g}$ associated with the Lie algebra $\mathfrak{g}$ (so it is logical to think of $\mathfrak{g}^{({\bf a})}$ as the vertex type). Ultimately, on Theorems 1.4 and 1.6, we say that an edge model of an electric Lie algebra of type $\mathfrak{g}$ is any subalgebra of $\mathfrak{g}$ isomorphic to $\mathfrak{g}^{({\bf b})}$.
Clearly, if all $b_{ij}=a_{ji}a_{i}a_{j}$ whenever $a_{ij}=-1$ in the notation of (1.2), then by Theorem 1.3, $\mathfrak{g}^{({\bf b})}\cong\mathfrak{g}^{({\bf a})}$. Note however, that the generalized electrical Lie algebra $sl_{n}^{({\bf b})}$ of the edge type $sl_{n}$ is also embedded into $sl_{n}$ for any ${\bf b}\in\mathbb{C}^{n}$ (here $b_{i}=b_{i,i+1}=b_{i+1,i}$ for $i=1,\ldots,n-2$).
Based on this, we can pose a natural
Problem 1.7.
Describe the set ${\mathcal{V}}_{A}$ of all tuples ${\bf b}=\{b_{ij}=b_{ji}|a_{ij}=-1\}$ such that $\mathfrak{g}^{(\bf b)}$ admits an edge model.
Theorem 1.3 gives an edge model for all ${\bf b}=(b_{ij})\in{\mathcal{V}}_{A}$ with $b_{ij}=a_{ji}a_{i}a_{j}$.
[…]
Theorem 1.8.
$sl_{n}^{(\bf b)}$ preserves the form $\omega_{\bf b}:=\sum\limits_{k=1}^{n-1}\left(\prod\limits_{i=k}^{n-2}(-b_{i})\right)v^{*}_{k}\wedge v^{*}_{k+1}$ in $\mathbb{C}^{n}=\mathbb{C}v_{1}\oplus\cdots\oplus\mathbb{C}v_{n}$. If $n$ is even then this form is non-degenerate as it follows from [11]. In this case $sl_{n}^{(\bf b)}$ is the intersection of $sp_{\omega_{\bf b}}\subset sl_{n}$ with the annihilator of $v^{1}=v_{1}-b_{1}v_{3}+b_{1}b_{3}v_{5}-b_{1}b_{3}b_{5}v_{7}+\cdots\in\mathbb{C}^{n}=V_{\omega_{1}}$ (where $\{v_{1},\ldots,v_{n}\}$ is the standard basis of $\mathbb{C}^{n}$), therefore, $sl_{n}^{(\bf b)}$ is isomorphic to $sp_{n-1}$ when $b_{1}\cdots b_{n-2}\neq 0$. If $n$ is odd, then the form $\omega_{\bf b}$ has a one-dimensional kernel that is an invariant of the action of $sl_{n}^{(\bf b)}$.
We prove Theorem 1.8 in Section 3.4. For $n=6$ the form $\Omega_{6}^{(\bf b)}$ is given in $v_{1},\ldots,v_{6}$ by its Gram matrix
$\Omega_{6}^{(\bf b)}=\begin{pmatrix}0&b_{1}b_{2}b_{3}b_{4}&0&0&0&0\\ -b_{1}b_{2}b_{3}b_{4}&0&-b_{2}b_{3}b_{4}&0&0&0\\ 0&b_{2}b_{3}b_{4}&0&b_{3}b_{4}&0&0\\ 0&0&-b_{3}b_{4}&0&-b_{4}&0\\ 0&0&0&b_{4}&0&1\\ 0&0&0&0&-1&0\end{pmatrix}.$
In view of Theorem 1.8, the electrical Lie algebra of type $sl_{n}$ has a trivial center if $n$ is odd and a one-dimensional center if $n$ is even. It is curious that the non-injective homomorphism $sl_{2k}^{(\bf b)}\cong sp_{2k-1}\to sl_{2k-1}$ is far from the natural embedding $sp_{2k-1}\subset sl_{2k-1}$.
We will use other principle unipotents for conjugation of the standard embedding into several new ones.
Example 1.9.
Conjugating with $g^{\prime}_{\bf a}=e^{a_{2}f_{2}}e^{a_{1}f_{1}}e^{a_{3}f_{3}}$ and $g^{\prime\prime}_{\bf a}=e^{a_{1}f_{1}}e^{a_{3}f_{3}}e^{a_{2}f_{2}}$ in $SL_{4}$ respectively, gives two new embeddings $sl_{4}^{(\bf a)}\hookrightarrow sl_{4}$:
$g^{\prime}_{\bf a}u_{i}{g^{\prime}}_{\bf a}^{-1}=e_{i}+\delta_{i,2}(b_{1}f_{1}+b_{3}f_{3}+b_{1}b_{3}[f_{1},[f_{3},f_{2}]])\ ,$
$g^{\prime\prime}_{\bf a}u_{i}{g^{\prime\prime}}_{\bf a}^{-1}=e_{i}+(1-\delta_{i,2})(b_{i}f_{2}+b[f_{i},f_{2}])$
for $i=1,2,3$, where we abbreviated $b_{1}=-a_{1}a_{2}$, $b_{3}=-a_{2}a_{3}$, $b=-a_{1}a_{2}a_{3}$.
We generalize the first embedding to all $\mathfrak{g}$ with a conical Dynkin diagram which includes all simply-laced $\mathfrak{g}$ (Theorem 2.1) and the second embedding to all $\mathfrak{g}$ with a star-like conical Dynkin diagram, including $D_{4}$ and $\hat{D}_{4}$ (Theorem 2.5). The following result gives such an edge model in types $B$ and $C$ and is of “conjugation type” as Theorem 1.6.
Theorem 1.10.
(a) Let $\mathfrak{g}=so_{2n+1}$ with $I=\{1,\ldots,n\}$ and the short root is $\alpha_{1}$. Then the assignments $u_{i}\mapsto e_{i}+b_{i-1,i}f_{i-1}$, $i=1,\ldots,n$ define an injective homomorphism of Lie algebras $so_{2n+1}^{(\bf b)}\hookrightarrow so_{2n+1}$ (with the convention $b_{01}=0$)
(b) Let $\mathfrak{g}=sp_{2n}$ with $I=\{1,\ldots,n\}$ and let the long root be $\alpha_{n}$. Then the assignments $u_{i}\mapsto e_{i}+b_{i-1,i}f_{i-1}-\delta_{i,n}\frac{1}{2}b_{n,n-1}^{2}[f_{n-1},[f_{n-1},f_{n}]]$ for $i=1,\ldots,n$ define an injective homomorphism of the Lie algebras $sp_{2n}^{(\bf b)}\hookrightarrow sp_{2n}$.
We prove Theorem 1.10 in Section 3.5 by using results of Section 2 where we also construct vertex models for electrical Lie algebras of several other Kac-Moody types, including $E_{6}$, $E_{7}$, $E_{8}$, $F_{4}$, $G_{2}$ and their affine versions as well (Theorems 2.1 and 2.3) along with some edge models (Theorem 2.9).
Our next result also gives an edge model for dihedral, type $D$, and affine $A$ types, however, unlike Theorem 1.10, we do not expect any intertwiner from the vertex model.
Theorem 1.11.
(a) Let $\mathfrak{g}=\mathfrak{g}_{A}$ be a Kac-Moody algebra of rank $2$ with $I=\{1,2\}$ such that $a_{21}\leq-2$. Then the assignments $u_{1}\mapsto e_{1}$, $u_{2}\mapsto e_{2}+b_{12}f_{1}$, define an injective homomorphism of Lie algebras $\mathfrak{g}^{(\bf b)}\hookrightarrow\mathfrak{g}$.
$b:=a_{21}a_{1}a_{2}$
(b) Let $\mathfrak{g}=so_{2n}$, $n\geq 3$ with $I=\{1,\ldots,n\}$ and with the branch at $i=n-2$. Then the assignments $u_{i}\mapsto\begin{cases}e_{i}+b_{i-1,i}f_{i-1}&\text{if $1\leq i\leq n-2$}\\ e_{i}+b_{n-2,i}f_{n-2}&\text{if $i\in\{n-1,n\}$}\\ \end{cases}$ $i=1,\ldots,n$, define an injective homomorphism of Lie algebras $so_{2n}^{(\bf b)}\hookrightarrow so_{2n}$.
(c) Let $\mathfrak{g}=\widehat{sl}_{n}$, the untwisted affine Lie algebra of type $\hat{A}_{n-1}$ with $I=\{0,\ldots,n-1\}$. Then the assignments $u_{i}\mapsto e_{i}+b_{i-1,i}f_{i-1}$, where $i-1$ is calculated modulo $n$ for $i\in I$, define an injective homomorphism of the Lie algebras $\widehat{sl}^{({\bf b})}\hookrightarrow\widehat{sl}_{n}$.
We prove Theorem 1.11 in Section 3.8.
Remark 1.12.
The affine electrical Lie algebra of type $\hat{A}_{n-1}$ from the theorem above was first discovered by T.Lam and A.Postnikov, see [7].
Theorem 1.13.
For any ${\bf b}\in\mathbb{C}^{n-1}$ one has:
(a) $sp^{({\bf b})}_{2n}\cong sl_{n}^{(b_{1},\ldots,b_{n-2})}\ltimes J$, where the first factor is a copy of $sl_{n}^{(b_{1},\ldots,b_{n-2})}$ in $sp^{({\bf b})}_{2n}$ generated by $u_{1},\ldots,u_{n-1}$ and $J$ is the Lie ideal of $sp^{({\bf b})}_{2n}$ generated by $u_{n}$ and $sl_{n}^{(b_{1},\ldots,b_{n-2})}$.
As $sl_{n}^{(b_{1},\ldots,b_{n-2})}$ module $J$ is isomorphic to $S^{2}V$ where $V$ is the restriction to $sl_{n}^{(b_{1},\ldots,b_{n-2})}\subset sl_{n}$ of the standard $sl_{n}$-module $V_{\omega_{1}}=\mathbb{C}^{n}$.
(b) If all $b_{i}\neq 0$ then, under then there is an isomorphism of Lie algebras $J\cong sl_{n+1}^{({\bf b^{\prime}})}$ for some choice of the parameters ${\bf b^{\prime}}$ (which we specify in the proof) and the Lie algebra $sp^{({\bf b})}_{2n}$ is naturally isomorphic to $sl_{n}^{(b_{1},\ldots,b_{n-2})}\ltimes sl_{n+1}^{({\bf b^{\prime}})}$.
We prove Theorem 1.13 in Section 3.6.
Remark 1.14.
It is well-known that $\mathfrak{h}\ltimes\mathfrak{g}\cong\mathfrak{h}\oplus\mathfrak{g}$ for any Lie algebra $\mathfrak{g}$ and it subalgebra $\mathfrak{h}$ under the right adjoint action of $\mathfrak{h}$ on $\mathfrak{g}$ (namely, the diagonal copy of $\mathfrak{h}$ in $\mathfrak{h}\ltimes\mathfrak{g}\cong$ commutes with $(0,\mathfrak{g})$).
Moreover, applying this to $\mathfrak{g}=sp_{n}$, $\mathfrak{h}=sp_{n-1}$, Theorems 1.8 and 1.13(b) produce, after an appropriate localization, an isomorphism
$sp^{({\bf b})}_{2n}\cong sl_{n}^{(b_{1},\ldots,b_{n-2})}\oplus sl_{n+1}^{(b_{1},\ldots b_{n-1})}$
recovering [10, Thm. 2.3.1] [^1] The electric algebras of $B_{n}$ type in [10] correspond to $C_{n}$ type in our approach..
We conclude with an illustration of Theorem 1.13 for $n=3$.
Example 1.15.
The Chevalley generators of $sp_{6}$ are given in terms of the folding of those for $sl_{6}$ under the flip $i\mapsto 6-i$:
$\tilde{e}_{1}=e_{1}+e_{5},\,\tilde{e}_{2}=e_{2}+e_{4},\,\tilde{e}_{3}=e_{3},\,\tilde{f}_{1}=f_{1}+f_{5},\,\tilde{f}_{2}=f_{2}+f_{4},\,\tilde{f}_{3}=f_{3}.$
The flip-invariant copy of $sl^{(\bf b)}_{6}$ is generated by the elements (Theorem 2.3 with $k=3$)
$\tilde{u}_{1}=e_{1},\,\tilde{u}_{2}=e_{2}+b_{1}f_{1},\,\tilde{u}_{3}=\tilde{e}_{3}+b_{2}\tilde{f}_{2}-b_{2}^{2}[\tilde{f}_{2},[\tilde{f}_{2},\tilde{f}_{3}]]/2$
$\tilde{u}_{4}=e_{4}+b_{1}f_{5},\,\tilde{u}_{5}=e_{5}$
the generators of ${sp}^{(\bf b)}_{6}$ are
$u_{1}=\tilde{u}_{1}+\tilde{u}_{5},\,u_{2}=\tilde{u}_{2}+\tilde{u}_{4},\,u_{3}=\tilde{u}_{3}.$
Introduce the elements of ${sp}^{(\bf b)}_{6}$
$w_{3}=u_{3},\,w_{2}=[u_{2},[u_{2},u_{3}]],\,w_{1}=[u_{1},[u_{1},w_{2}]].$
These elements on the one hand generate the ideal $J$ and on the other hand a copy of $sl_{4}^{({\bf b^{\prime}})}$ where ${\bf b^{\prime}}=\{-32b^{2}_{1}b^{2}_{2}\,,-8b^{2}_{2}\}$. This shows a splitting of $sp_{6}^{({\bf b})}$ into a semidirect product as claimed in the statement of the theorem.
Moreover, as Remark 1.14 claims there is copy of $sl_{3}^{({\bf b^{\prime}})\backslash-8b^{2}_{2}}$ that commutes with the above copy of $sl_{4}^{({\bf b^{\prime}})}$. It is generated by the elements
$v_{1}=-8b_{1}b_{2}u_{1}+w_{1},\,v_{2}=4b_{2}u_{2}+w_{2}$
After the appropriate localization
$sp^{(\bf b)}_{6}\cong sl_{2}^{(b_{1})}\oplus sl_{3}^{({b_{1},b_{2}})}$
where the generators of the summands are
$\{u_{1}-\frac{w_{1}}{8b_{1}b_{2}},\,u_{2}+\frac{w_{2}}{4b_{2}}\}\,\text{and}\,\{w_{3},\,-\frac{w_{2}}{8b_{2}},\,\frac{w_{1}}{4b_{1}b_{2}}\}$
respectively.
2. Other vertex and edge models of electrical Lie algebras
In this section, we generalize Theorem 1.4, Example 1.9 and Theorems 1.10, 1.11 to other semisimple and Kac-Moody algebras.
To any (generalized) Cartan matrix $A=(a_{ij},i,j\in I)$ we assign its graph $\Gamma(A)$ on the vertex set $I$ such that $(ij)$ is an edge iff $a_{ij}<0$.
We say that a rooted tree is conical if the only possible branch is at the root (which we always denote by 00).
For example, a linear graph with the set of vertex $[-m,n]$, $m,n\geq 0$ is a conical tree. Also, all Dynkin diagrams of types $ADE$ are conical rooted trees.
Any rooted tree $\Gamma$ with root 00 is naturally a layered partial order $\prec$ on $\Gamma$ with the maximal element 00 and whose minimal elements are the leaves. Furthermore, for any $i\in\Gamma\setminus\{0\}$ denote by $i^{+}\in I$ the parent of $i$ in $\prec$, that is, the smallest element $j$ such that $i\prec j$.
Likewise, if $\Gamma$ is a conical tree, for any non-leaf $i\in\Gamma\setminus\{0\}$ denote by $i^{-}\in I$ the only son of $i$ in $\prec$.
Theorem 2.1.
Suppose that $\Gamma(A)$ is a conical tree with the root 00 and $a_{ij}\in\{0,-1\}$ for all $i,j\in I$ such that $j\neq 0$. Then $(Ad~{}g_{\bf a})(u_{i})=$
$\begin{cases}e_{i}+b_{i^{-}}f_{i^{-}}&\text{if $i\neq 0$ }\\ e_{0}-a_{0}{\bf f}-\frac{a_{0}^{2}}{2}[{\bf f},[{\bf f},f_{0}]]-\sum\limits_{k\geq 3}\frac{a_{0}^{2}}{k!}(ad~{}{\bf f}+a_{0}[f_{0},{\bf f}])^{k-1}([{\bf f},[{\bf f},f_{0}]])&\text{if $i=0$}\end{cases}$
for $i\in\Gamma(A)$, where we abbreviated $b_{i^{-}}:=-a_{i}a_{i^{+}}$ for $i\in I$ (with the convention $b_{i^{-}}=0$ for any leaf $i$ of $\Gamma$), ${\bf f}:=\sum\limits_{j\in I:j^{+}=0}a_{j}f_{j}$ and $g_{\bf a}=e^{a_{0}f_{0}}\prod\limits_{i\neq 0}e^{a_{i}f_{i}}$, where the product is decreasing (i.e., $i^{+}$ always precedes $i$).
Proof.
We need the following
Proposition 2.2.
Suppose that $\Gamma(A)$ is a conical tree with the root 00 and such that $a_{ij}\in\{0,-1\}$ for all $i,j\in I$ such that $0\notin\{i,j\}$. Then
$(Ad~{}g_{\bf a})(u_{i})=\begin{cases}e_{i}+b_{i^{-}}f_{i^{-}}&\text{if $i\neq 0$ }\\ e_{0}+a_{0}[\hat{\bf f},h_{0}+a_{0}f_{0}]-a_{0}^{2}\sum\limits_{k\geq 2}\frac{1}{k!}(ad~{}\hat{\bf f})^{k}(f_{0})&\text{if $i=0$}\end{cases}$
for $i\in\Gamma(A)$, where we abbreviated $b_{i^{-}}:=-a_{i}a_{i^{+}}$ for $i\in I$ (with the convention $b_{i^{-}}=0$ for any leaf $i$), ${\bf f}:=\sum\limits_{j\in I:j^{+}=0}a_{j}f_{j}$ and $g_{\bf a}=e^{a_{0}f_{0}}\prod\limits_{i\neq 0}e^{a_{i}f_{i}}$, where the product is decreasing (i.e., $i^{+}$ always precedes $i$), and $\hat{\bf f}=(Ad~{}e^{a_{0}f_{0}})({\bf f})=\sum\limits_{k\geq 0}\frac{a_{0}^{k}}{k!}(ad~{}f_{0})^{k}({\bf f})$.
Proof.
Indeed, if $i\neq 0$, then $g_{\bf a}=g^{\prime}g^{\prime\prime}$ where $g^{\prime}=e^{a_{0}f_{0}}\prod\limits_{i\in I^{\prime}}e^{a_{i}f_{i}}$ and $g^{\prime\prime}=\prod\limits_{j\in I^{\prime\prime}}e^{a_{j}f_{j}}$, both products decreasing where $I^{\prime\prime}=\{j:j\preceq i\}$, $I^{\prime}=(I\setminus\{0\})\setminus I^{\prime\prime}$.
As in the proof of Theorem 1.4, $(Ad~{}g^{\prime\prime})(u_{i})=e_{i}+b_{i^{-}}f_{i^{-}}$ hence
$(Ad~{}g_{\bf a})(u_{i})=(Ad~{}g^{\prime})(e_{i}+b_{i^{-}}f_{i^{-}})=e_{i}+b_{i^{-}}f_{i^{-}}$
because both $e_{i}$ and $f_{i^{-}}$ are fixed by $Ad~{}g^{\prime}$.
This proves the first case.
Now let $i=0$. Then
$(Ad~{}g_{\bf a})(u_{0})=(Ad~{}e^{a_{0}f_{0}}\prod\limits_{i\neq 0}e^{a_{i}f_{i}})(u_{0})$
$=(Ad~{}e^{a_{0}f_{0}}\prod\limits_{i\in I:i^{+}=0}e^{a_{i}f_{i}})(u_{0})=(Ad~{}e^{a_{0}f_{0}}e^{\bf f})(u_{0})$
because $[f_{i},f_{j}]=0$ if $i^{+}=j^{+}=0$. Furthermore,
$(Ad~{}e^{\bf f})(u_{0})=(Ad~{}e^{\bf f})(e_{0}+a_{0}h_{0}-a_{0}^{2}f_{0})=e_{0}+a_{0}(Ad~{}e^{\bf f})(h_{0})-a_{0}^{2}(Ad~{}e^{\bf f})(f_{0})$
$=u_{0}+a_{0}[{\bf f},h_{0}]-a_{0}^{2}((Ad~{}e^{\bf f})(f_{0})-f_{0})$
because $[{\bf f},h_{0}]=\sum\limits_{i\in I:i^{+}=0}a_{i}a_{0i}f_{i}$. Therefore, $[{\bf f},[{\bf f},h_{0}]]=0$. Finally,
$(Ad~{}g_{\bf a})(u_{0})=(Ad~{}e^{a_{0}f_{0}})(u_{0}+a_{0}[{\bf f},h_{0}]-a_{0}^{2}((Ad~{}e^{\bf f})(f_{0})-f_{0}))$
$=e_{0}+a_{0}[\hat{\bf f},h_{0}+2a_{0}f_{0}]-a_{0}^{2}((Ad~{}e^{\hat{\bf f}})(f_{0})-f_{0})$
because $(Ad~{}e^{a_{0}f_{0}})(h_{0})=h_{0}+2a_{0}f_{0}$.
This proves the second case. The proposition is proved. $\square$
Since $a_{0i}=-1$ whenever $i^{+}=0$ hence $[f_{0},[f_{0},{\bf f}]]=0$, we obtain
$\hat{\bf f}={\bf f}+a_{0}[f_{0},{\bf f}]\ .$
Also, $[{\bf f},h_{0}]=-{\bf f}$ and $[[f_{0},{\bf f}],h_{0}]=[f_{0},{\bf f}]$, therefore,
$[\hat{\bf f},h_{0}+a_{0}f_{0}]=[{\bf f}+a_{0}[f_{0},{\bf f}],h_{0}+a_{0}f_{0}]=-{\bf f}\ .$
Finally, note that
$[\hat{\bf f},f_{0}]=[{\bf f}+a_{0}[f_{0},{\bf f}],f_{0}]=[{\bf f},f_{0}]$
and
$[\hat{\bf f},[{\bf f},f_{0}]]=[{\bf f}+a_{0}[f_{0},{\bf f}],[{\bf f},f_{0}]]]=[{\bf f},[{\bf f},f_{0}]]]$
The theorem is proved. $\square$
Theorem 2.1 covers the cases when $\mathfrak{g}$ is of types $E_{6},E_{7},E_{8}$, and $F_{4}$.
Viewing the Dynkin diagram of $sl_{n}$ as a conical tree with a “root” $0=k\in\{1,\ldots,n-1\}$ we obtain the following corollary of Theorem 2.1, which generalizes Theorem 1.4.
Theorem 2.3.
For any $k\in\{1,\ldots,n-1\}$ the assignments
$u_{i}\mapsto\begin{cases}e_{i}+b_{i-1}f_{i-1}&\text{if $i<k$}\\ e_{i}+b_{i}f_{i+1}&\text{if $i>k$}\\ e_{k}+b_{k-1}f_{k-1}+b_{k}f_{k+1}+b_{k-1}b_{k}[f_{k-1},[f_{k+1},f_{k}]]&\text{if $i=k$}\end{cases}$
define an injective homomorphism $sl_{n}^{({\bf b})}\hookrightarrow sl_{n}$ (where $b_{i}:=-a_{i}a_{i+1}$).
We prove Theorem 2.3 in Section 3.5.
The following is an immediate corollary of Theorem 2.1.
Corollary 2.4.
Let $\Gamma(A)$ be a conical star tree (i.e., every non-root is a leaf) with $a_{i0}=a_{0i}=-1$ for all $i\in I\setminus\{0\}$. Then the assignments
$u_{i}\mapsto e_{0}-a_{0}\delta_{i,0}({\bf f}-\frac{a_{0}}{2}[{\bf f},[{\bf f},f_{0}]]-a_{0}\sum\limits_{k\geq 3}\frac{1}{k!}(ad~{}{\bf f}+a_{0}[f_{0},{\bf f}])^{k-1}([{\bf f},[{\bf f},f_{0}]]))\ ,$
where ${\bf f}:=\sum\limits_{j\neq 0}a_{j}f_{j}$, define an injective homomorphism $\mathfrak{g}^{({\bf a})}\hookrightarrow\mathfrak{g}$.
Theorem 2.5.
Suppose that $\Gamma(A)$ is a conical tree with the root 00, $J_{+}$ is a set of leaves of $\Gamma(A)$ attached to the 00, and $a_{ij}\in\{0,-1\}$ for all $i,j\in I\setminus J_{+}$ such that $j\neq 0$ and (i.e., $J_{+}^{+}=\{0\}$) such that $a_{i0}=-1$ for all $i\in J_{+}$. Then
$(Ad~{}g_{\bf a})(u_{i})=\begin{cases}e_{i}+b_{i}f^{\prime}_{i}-\frac{b_{i}^{2}}{2}[[f_{i},f^{\prime}_{i}],f^{\prime}_{i}]\\ -a_{i}^{2}\sum\limits_{k\geq 3}\frac{a_{0}^{k}}{k!}(ad~{}(f^{\prime}_{i}+a_{i}[f_{i},f^{\prime}_{i}]))^{k-2}([[f_{i},f^{\prime}_{i}],f^{\prime}_{i}])&\text{if $i\in J_{+}$}\\ e_{0}-a_{0}{\bf f}-\frac{a_{0}^{2}}{2}[{\bf f},[{\bf f},f^{\prime}_{0}]]\\ -a_{0}^{2}\sum\limits_{k\geq 3}\frac{1}{k!}((ad~{}{\bf f}+a_{0}[f^{\prime}_{0},{\bf f}])^{k-1}([{\bf f},[{\bf f},f^{\prime}_{0}]]))&\text{if $i=0$}\\ e_{i}+b_{i^{-}}f_{i^{-}}&\text{otherwise}\\ \end{cases}$
for $i\in\Gamma(A)$, where we abbreviated $b_{i^{-}}:=-a_{i}a_{i^{+}}$ for $i\in I\setminus J_{+}$, $b_{i}:=-a_{0}a_{i}$, $i\in J_{+}$ (with the convention $b_{i^{-}}=0$ for any leaf $i$), ${\bf f}_{+}:=\sum\limits_{j\in J_{+}}a_{j}f_{j}$, ${\bf f}:=\sum\limits_{j\in I\setminus J_{+}:j^{+}=0}a_{j}f_{j}$, $f^{\prime}_{0}:=\sum\limits_{k\geq 0}\frac{1}{k!}(ad~{}{\bf f}_{+})^{k}(f_{0})$, $f^{\prime}_{i}:=\sum\limits_{k\geq 0}\frac{1}{k!}(ad~{}({\bf f}_{+}-a_{i}f_{i}))^{k}(f_{0})$, $i\in J_{+}$, and $g_{\bf a}=e^{{\bf f}_{+}}e^{a_{0}f_{0}}\prod\limits_{i\in I\setminus(\{0\}\cup J_{+})}e^{a_{i}f_{i}}$, where the product is decreasing (i.e., $i^{+}$ always precedes $i$).
Proof.
The case $i\neq 0$, $i\notin J_{+}$ follows from the proof of Theorem 2.1.
Now let $i\in J_{+}$. Note that $(Ad~{}e^{a_{j}f_{j}})(u_{i})=u_{i}$ for all $j\neq 0$, $j\notin J_{+}$. Also, $i\in J_{+}$:
$(Ad~{}e^{a_{i}f_{i}}e^{a_{0}f_{0}})(u_{i})=(Ad~{}e^{a_{i}f_{i}}e^{a_{0}f_{0}}e^{-a_{i}f_{i}})(e_{i})=(Ad~{}e^{a_{0}(f_{0}+a_{i}[f_{i},f_{0}])})(e_{i})$
Taking into account that
$(ad~{}((f_{0}+a_{i}[f_{i},f_{0}]))(e_{i})=a_{i}[[f_{i},f_{0}],e_{i}]$
$=a_{i}[[f_{i},e_{i}],f_{0}]=-a_{0}a_{i}[h_{i},f_{0}]=-a_{i}f_{0}$
and
$(ad~{}(f_{0}+a_{i}[f_{i},f_{0}]))^{k-1}(f_{0})=a_{i}[[f_{i},f_{0}],f_{0}]]\ ,$
we obtain
$(Ad~{}e^{a_{i}f_{i}}e^{a_{0}f_{0}})(u_{i})=e_{i}-a_{0}a_{i}f_{0}+\sum_{k\geq 2}\frac{a_{0}^{k-1}}{k!}(ad~{}(f_{0}+a_{i}[f_{i},f_{0}]))^{k-1}(a_{i0}a_{0}a_{i}f_{0})$
$=e_{i}+b_{i}f_{0}-\frac{b_{i}^{2}}{2}[[f_{i},f_{0}],f_{0}]-a_{0}a_{i}^{2}\sum_{k\geq 3}\frac{a_{0}^{k-1}}{k!}(ad~{}(f_{0}+a_{i}[f_{i},f_{0}]))^{k-2}([[f_{i},f_{0}],f_{0}])\ .$
Finally,
$Ad~{}g_{\bf a}(u_{i})=(Ad~{}e^{{\bf f}_{+}}e^{a_{0}f_{0}})(u_{i})=(Ad~{}e^{{\bf f}_{+}-a_{i}f_{i}})((Ad~{}e^{a_{i}f_{i}}e^{a_{0}f_{0}})(u_{i}))$
$e_{i}+b_{i}f^{\prime}_{0i}-\frac{b_{i}^{2}}{2}[[f_{i},f^{\prime}_{0i}],f^{\prime}_{0i}]-a_{i}^{2}\sum_{k\geq 3}\frac{a_{0}^{k}}{k!}(ad~{}(f^{\prime}_{0i}+a_{i}[f_{i},f^{\prime}_{0i}]))^{k-2}([[f_{i},f^{\prime}_{0i}],f^{\prime}_{0i}])\ .$
Now consider the remaining case $i=0$. Then
$Ad~{}g_{\bf a}(u_{0})=(Ad~{}e^{{\bf f}_{+}}e^{a_{0}f_{0}}e^{\bf f}e^{-a_{0}f_{0}})(e_{0})=(Ad~{}e^{{\bf f}_{+}}e^{\hat{\bf f}})(e_{0})=Ad~{}e^{{\bf f}_{+}}(e^{\bf f}(e_{0}))$
where $\hat{\bf f}=(Ad~{}e^{a_{0}f_{0}})({\bf f})={\bf f}+a_{0}[f_{0},{\bf f}]$.
Finally, using the argument from the proof of Theorem 2.1 for $i=0$ and replacing $f_{0}$ with $f^{\prime}_{0}=Ad~{}e^{{\bf f}_{+}}(f_{0})$, we finish the proof of Theorem 2.5.
$\square$
The following is an immediate corollary of Theorem 2.5.
Corollary 2.6.
In the assumptions of Theorem 2.5 assume additionally, that $I\setminus J_{+}$ is of type $A$. Then
(a) the assignments
$u_{i}\mapsto\begin{cases}e_{i}+b_{i^{-}}f_{i^{-}}&\text{if $i\notin J_{+}$ }\\ e_{i}+b_{i}f_{0}+b_{i}\sum\limits_{k\geq 1}\frac{1}{k!}(ad~{}({\bf f}_{+}-a_{i}f_{i}))^{k}(f_{0})&\text{if $i\in J_{+}$}\\ \end{cases}$
define an injective homomorphism $\mathfrak{g}^{({\bf a})}\hookrightarrow\mathfrak{g}$.
(b) Suppose that $I=\{1,\ldots,n\}$, $\{1,\ldots,n-1\}$ is of type $A_{n-1}$, and $a_{in}=a_{ni}=\delta_{ik}$. Then the assignments
$u_{i}\mapsto\begin{cases}e_{i}+b_{i+1}f_{i+1}&\text{if $1\leq i<k$}\\ e_{k}+b_{k-1,k}f_{k-1}+b_{k,k+1}f_{k+1}\\ +b_{k-1,k}b_{k,k+1}[f_{k-1},[f_{k+1},f_{k}+a_{n}[f_{n},f_{k}]]]&\text{if $i=k$}\\ \end{cases}$
where $b_{ij}:=a_{ij}a_{i}a_{j}$, define an injective homomorphism $\mathfrak{g}^{({\bf a})}\hookrightarrow\mathfrak{g}$.
Theorem 2.7.
Let $\mathfrak{g}=\mathfrak{g}_{A}$ be a Kac-Moody algebra with $I=\{1,\ldots,r\}$ such that the Cartan matrix $A$ of $\mathfrak{g}$ satisfies $a_{ij}a_{ji}\neq 0$ iff $|i-j|=1$ and $a_{ij}\leq a_{i-1,j-1}$ for all distinct $i,j=2,\ldots,r$.
Suppose also that the following three conditions hold.
$\bullet$ Either $a_{12}=a_{21}=-1$ or $a_{21}<-1$.
$\bullet$ Either $a_{i-1,i}a_{i,i+1}>1$ or $a_{i-1,i}=a_{i,i+1}=a_{i,i-1}=-1$ for $i=2,\ldots,r$.
$\bullet$ Either $a_{i,i-1}a_{i+1,i}>1$ or $a_{i+1,i}=a_{i,i-1}=a_{i,i+1}=-1$ for $i=2,\ldots,r$.
Then the assignments $u_{i}\mapsto\begin{cases}e_{1}&\text{if $i=1$}\\ e_{i}+b_{i-1}f_{i-1}&\text{if $2\leq i\leq r$}\\ \end{cases}$ define a homomorphism of Lie algebras ${\mathfrak{g}^{\prime}}^{(\bf b)}\to\mathfrak{g}$, where we abbreviated $b_{j}:=a_{j+1,j}a_{j}a_{j+1}$ for $j=1,\ldots,r-1$, similarly to Theorem 1.10 and $A^{\prime}$ is $I\times I$ the Cartan matrix of $\mathfrak{g}^{\prime}$ given by $a^{\prime}_{12}=a_{12}$ $a^{\prime}_{21}=a_{21}$ and $a^{\prime}_{ij}=\min(a_{i-1,j-1},a_{ij})$ for $i,j=2,\ldots,r$
Theorem 2.8.
Suppose that $I=\{1,\ldots,r+d\}$, $r\geq 1$, $d\geq 1$ such that $a_{ij}a_{ji}\neq 0$ for distinct $i,j$ iff either $i,j\leq r$ and $|i-j|=1$ or $\min(i,j)=r$, the restriction of $A$ to $\{1,\ldots,r,r+i\}$ satisfies the assumptions of Theorem 2.7 for $i=1,\ldots,d$ and $a_{r+i,r+j}=0$ for all distinct $i,j=1,\ldots,d$. Then the assignments
$u_{i}\mapsto\begin{cases}e_{1}&\text{if $i=1$}\\ e_{i}+b_{i-1}f_{i-1}&\text{if $2\leq i\leq r$}\\ e_{i}+b_{i}f_{r}&\text{if $r<i\leq r+d$}\\ \end{cases}$
define a homomorphism of Lie algebras ${\mathfrak{g}^{\prime\prime}}^{(\bf b)}\to\mathfrak{g}$, where we abbreviated $b_{i}:=a_{ir}a_{r}a_{i}$ for $i=r+1,\ldots,r+d$ and $A^{\prime\prime}$ is the $I\times I$ Cartan matrix of $\mathfrak{g}^{\prime\prime}$ with $a^{\prime\prime}_{ij}=a^{\prime}_{ij}$ whenever $i,j=1,\ldots,r$, $a^{\prime\prime}_{ij}=\min(a_{i-1,j-1},a_{ij})$ whenever $\min(i,j)=r$, $i\neq j$ and the remaining $a^{\prime\prime}_{ij}=0$.
We prove Theorems 2.7 and 2.8 in Section 3.7.
Taking $\mathfrak{g}^{\prime\prime}=\mathfrak{g}$ in Theorem 2.8, we obtain the following
Theorem 2.9.
Suppose that $I=\{1,\ldots,r+d\}$, $a_{ij}a_{ji}\neq 0$ for distinct $i,j$ iff either $i,j\leq r$ and $|i-j|=1$ or $\min(i,j)=r$ (as in Theorem 2.8) and the Cartan matrix of $\mathfrak{g}$ satisfies
$\bullet$ Either $a_{12}=a_{21}=-1$ or $a_{21}<-1$.
$\bullet$ Either $a_{i-1,i}=a_{i,i+1}=a_{i,i-1}=-1$ or $a_{i,i+1}<-1$ and $a_{i,i+1}\leq a_{i-1,i}$ for $i=2,\ldots,r$
$\bullet$ Either $a_{i+1,i}=a_{i,i-1}=a_{i,i+1}=-1$ or $a_{i+1,i}<-1$ and $a_{i+1,i}\leq a_{i,i-1}$ for $i=2,\ldots,r$.
$\bullet$ $a_{ij}\leq a_{i-1,j-1}$ whenever $\min(i,j)=r$, $i\neq j$.
Then the subalgebra generated by $u_{i}:=\begin{cases}e_{1}&\text{if $i=1$}\\ e_{i}+b_{i-1}f_{i-1}&\text{if $2\leq i\leq r$}\\ e_{i}+b_{i}f_{r}&\text{if $r<i\leq r+d$}\\ \end{cases}$ is naturally isomorphic to $\mathfrak{g}^{(\bf b)}$ for any ${\bf b}=(b_{1},\ldots,b_{n-1})$.
We prove Theorem 2.9 in Section 3.8.
As in the proof of Theorem 1.3, all results of this section are valid if one replaces $\mathbb{C}$ with and $\mathbb{C}$-algebra containing all $a_{i}$ (resp. all $b_{ij}$).
3. Proof of main results
3.1. Proof of Theorem1.2
Let $e_{i^{r}j}:=\frac{1}{r!}(ad~{}e_{i})^{r}(e_{j})$, $f_{i^{r}j}:=\frac{1}{r!}(ad~{}f_{i})^{r}(f_{j})$, and $u_{i^{r}j}:=\sum\limits_{k=0}^{r}a_{i}^{r-k}\binom{a_{ij}+r-1}{r-k}(e_{i^{k}j}-(-1)^{r}a_{i}^{2k}a_{j}^{2}f_{i^{k}j})$ for any distinct $i,j\in I$ and $r\geq 1$.
We need the following result.
Proposition 3.1.
$[u_{i},u_{i^{r}j}]=(r+1)u_{i^{r+1}j}$ for all distinct $i,j\in I$ and $r\geq 1$.
Proof.
The following is immediate.
Lemma 3.2.
$[f_{i},e_{i^{k}j}]=-(a_{ij}+k-1)e_{i^{k-1}j}$, $[e_{i},f_{i^{k}j}]=-(a_{ij}+k-1)f_{i^{k-1}j}$ for all $k\geq 0$ (with the convention $f_{i^{-1}j}=e_{i^{-1}j}=0$).
Thus, we have for $r\geq 1$:
$[e_{i},u_{i^{r}j}]=\sum_{k=0}^{r}a_{i}^{r-k}\binom{a_{ij}+r-1}{r-k}([e_{i},e_{i^{k}j}]-(-1)^{r}a_{i}^{2k}a_{j}^{2}[e_{i},f_{i^{k}j}])$
$=\sum_{k=0}^{r}a_{i}^{r-k}\binom{a_{ij}+r-1}{r-k}((k+1)e_{i^{k+1}j}+(-1)^{r}(a_{ij}+k-1)a_{i}^{2k}a_{j}^{2}f_{i^{k-1}j})\ ,$
$[h_{i},u_{i^{r}j}]=\sum_{k=0}^{r}a_{i}^{r-k}\binom{a_{ij}+r-1}{r-k}([h_{i},e_{i^{k}j}]-(-1)^{r}a_{i}^{2k}a_{j}^{2}[h_{i},f_{i^{k}j}])$
$=\sum_{k=0}^{r}a_{i}^{r-k}\binom{a_{ij}+r-1}{r-k}(2k+a_{ij})(e_{i^{k}j}+(-1)^{r}a_{i}^{2k}a_{j}^{2}f_{i^{k}j})\ ,$
$[f_{i},u_{i^{r}j}]=\sum_{k=0}^{r}a_{i}^{r-k}\binom{a_{ij}+r-1}{r-k}([f_{i},e_{i^{k}j}]-(-1)^{r}a_{i}^{2k}a_{j}^{2}[f_{i},f_{i^{k}j}])$
$=\sum_{k=0}^{r}a_{i}^{r-k}\binom{a_{ij}+r-1}{r-k}(-(a_{ij}+k-1)e_{i^{k-1}j}-(-1)^{r}(k+1)a_{i}^{2k}a_{j}^{2}f_{i^{k+1}j})$
Therefore,
$[u_{i},u_{i^{r}j}]=\sum_{k=0}^{r+1}c_{k}a_{i}^{r+1-k}(e_{i^{k}j}+(-1)^{r}a_{i}^{2k}a_{j}^{2}f_{i^{k}j})$
where $c_{r+1}=r+1$ and
$c_{k}=(r-y)\binom{x}{y+1}+(x+r+1-2y)\binom{x}{y}+(x-y+1)\binom{x}{y-1}$
$=(r+1)\binom{x+1}{y+1}-(y+1)\binom{x}{y+1}+(x-2y)\binom{x}{y}+y\binom{x}{y}=(r+1)\binom{x+1}{y+1}$
because $\binom{x}{y+1}+\binom{x}{y}=\binom{x+1}{y+1}$, $(y+1)\binom{x}{y+1}=(x-y)\binom{x}{y}$, $(x-y+1)\binom{x}{y-1}=y\binom{x}{y}$, where we abbreviated $x:=a_{ij}+r-1$, $y:=r-k$.
This proves Proposition 3.1. $\square$
Furthermore,
$[u_{i},u_{j}]=[e_{i}+a_{i}h_{i}-a_{i}^{2}f_{i},e_{j}+a_{j}h_{j}-a_{j}^{2}f_{j}]=e_{ij}+a_{i}^{2}a_{j}^{2}f_{ij}$
$-a_{j}a_{ji}e_{i}+a_{i}a_{ij}e_{j}+a_{j}^{2}a_{i}a_{ij}f_{j}-a_{i}^{2}a_{j}a_{ji}f_{i}=u_{ij}-a_{j}a_{ji}(e_{i}+a_{i}^{2}f_{i})$
Taking into account that
$[u_{i},e_{i}+a_{i}^{2}f_{i}]=[e_{i}+a_{i}h_{i}-a_{i}^{2}f_{i},e_{i}+a_{i}^{2}f_{i}]=2a_{i}u_{i}$
and using Proposition 3.1 with $r=2$, we obtain
$(ad~{}u_{i})^{2}(u_{j})=2u_{i^{2}j}-2a_{i}a_{j}a_{ji}u_{i}\ .$
In turn, this and Proposition 3.1 imply by an induction in $r\geq 3$ that
$(ad~{}u_{i})^{r}(u_{j})=r!u_{i^{r}j}$
for all $r\geq 3$.
Finally, if $r=1-a_{ij}$ then $\binom{a_{ij}+r-1}{r-k}=0$ for all $k<r$ and $e_{i^{r}j}=f_{i^{r}j}=0$ which implies that $u_{i^{r}j}=0$.
Theorem 1.2 is proved. $\square$
3.2. Proof of Theorem 1.3
We apply the results of Appendix to our situation as follows.
Let $R$ be any $\mathbb{C}$-algebra, fix a family ${\bf a}=(a_{i},i\in I)$ of elements of $R$.
Denote $A:=R\otimes U(\mathfrak{g})$. By definition, the subalgebra $B$ of $A$ generated by $X:=\{u_{i}=e_{i}+a_{i}h_{i}-a_{i}^{2}f_{i},i\in I\}$ over $R$ is $U(\mathfrak{g}^{(\bf a)})$. According to (1.2), the kernel of the canonical homomorphism $\varphi_{X,B}:R\langle X\rangle\twoheadrightarrow B$ contains the set $Y:=\{(ad~{}u_{i})^{1-a_{ij}}(u_{j})+2\delta_{a_{ij},-1}a_{ji}a_{i}a_{j}u_{i}:i,j\in I,i\neq j\}$.
Let us define a filtration on $A$ by setting $\deg e_{i}=1$, $\deg f_{i}=\deg h_{i}=0$ for $i\in I$. Clearly, $grA$ is generated by $e_{i},f_{i},h_{i}$ subject to the usual Serre relations except for $[e_{i},f_{i}]=h_{i}$ which is replaced by $[e_{i},f_{i}]=0$. In particular, the subalgebra $B_{0}$ of $grA$ generated by $e_{i},i\in I$ is naturally isomorphic to $R\otimes U(\mathfrak{n})$.
Then the natural inclusion of $R$-algebras $B\subset A$ is compatible with the filtration and defines a natural inclusion of graded algebras $grB\subset B_{0}$ because $gr~{}u_{i}=e_{i}$ for $i\in I$ (that is, $grX=\{e_{i},i\in I\}$).
Clearly, $gr~{}Y=\{(ad~{}e_{i})^{1-a_{ij}}(e_{j}):i,j\in I,i\neq j\}$ and is the presentation of $B_{0}$, i.e., $gr~{}Y$ generates the kernel of the canonical homomorphism $\varphi_{grX,grB}:R\langle grX\rangle\twoheadrightarrow grB=B_{0}$.
Thus, all conditions of Proposition 4.1 are met. Thus, the proposition guarantees that the relations (1.2) give a presentation of $B=U(\mathfrak{g}^{(\bf a)})$ as an associative algebra over $R$ hence of the Lie algebra $\mathfrak{g}^{(\bf a)}$ over $R$.
The theorem is proved. $\square$
3.3. Proof of Theorem 1.4
Clearly, $u_{i}=(Ad~{}e^{-a_{i}f_{i}})(e_{i})$ for all $i\in I=\{1,\ldots,n-1\}$.
Then for $i=1$ we have
$(Ad~{}g_{\bf a})(u_{1})=(Ad~{}e^{a_{n-1}f_{n-1}}\cdots e^{a_{2}f_{2}})(e_{1})=e_{1}\ .$
Now let $i\geq 2$. Then
$(Ad~{}e^{a_{i}f_{i}}\cdots e^{a_{2}f_{2}}e^{a_{1}f_{1}})(u_{i})=(Ad~{}e^{a_{i}f_{i}}e^{a_{i-1}f_{i-1}}e^{-a_{i}f_{i}})(e_{i})$
$=(Ad~{}e^{a_{i-1}(f_{i-1}+a_{i}[f_{i},f_{i-1}])})(e_{i})=(Ad~{}e^{a_{i-1}a_{i}[f_{i},f_{i-1}]})(e_{i})$
$=e_{i}+a_{i-1}a_{i}[[f_{i},f_{i-1}],e_{i}]=e_{i}-a_{i-1}a_{i}f_{i-1}\ .$
Here we used the facts that $(Ad~{}e^{af_{j}})(e_{i})=e_{i}$ if $i\neq j$, $(Ad~{}e^{af_{j}})(f_{i})=f_{i}$ whenever $|i-j|>1$ and $(Ad~{}e^{a_{i}f_{i}})(f_{i-1})=f_{i-1}+a_{i}[f_{i},f_{i-1}]$ because $a_{i,i-1}=-1$. This is based on the following general fact
$(Ad~{}e^{x})(y)=e^{ad~{}x}(y)\ .$ (3.4)
Finally,
$(Ad~{}g_{\bf a})(u_{i})=(Ad~{}e^{a_{n-1}f_{n-1}}\cdots e^{a_{i+1}f_{i+1}})(e_{i}-a_{i-1}a_{i}f_{i-1})=e_{i}-a_{i-1}a_{i}f_{i-1}\ .$
The theorem is proved. $\square$
3.4. Proof of Theorem 1.8
Recall that $sl_{n}$ acts on $V^{*}$ via
$e_{i}(v_{j}^{*})=\delta_{ij}v^{*}_{i+1},~{}f_{i}(v_{j}^{*})=\delta_{i+1,j}v^{*}_{i}\,$
where $\{v^{*}_{i}\}$ is the basis of $V^{*}$ dual to $\{v_{i}\}$. Therefore, a direct calculation shows that the generators $u_{i}=e_{i}+b_{i-1}f_{i-1}$ for $i=1,\ldots,n-1$ of $sl_{n}^{({\bf b})}$ preserve
$\omega_{\bf b}:=\sum\limits_{k=1}^{n-1}\left(\prod\limits_{i=k}^{n-2}(-b_{i})\right)v^{*}_{k}\wedge v^{*}_{k+1}$ (3.5)
It is immediate that $v^{1}\in V$ is invariant under the restriction from $sl_{n}$-action on $V:=V_{\omega_{1}}$ to the $sl_{n}^{({\bf b})}$-action on $V$. This, in particular, implies that the orthogonal complement $V^{1}\subset V^{*}$ to $v^{1}$ is a (co-dimension $1$) $sl_{n}^{({\bf b})}$-submodule of $V^{*}$.
It is easy to see that $\omega_{\bf b}\in\Lambda^{2}V^{*}$ in fact belongs to $\Lambda^{2}V^{1}$ for odd $n$. Clearly, in this case vectors $w_{i}=\begin{cases}b_{i}v_{i}^{*}+v_{i+2}^{*}&\text{if $i$ is odd}\\ v_{i}^{*}&\text{if $i$ is even}\\ \end{cases}$ $i=1,\ldots,n-1$ form a basis of $V^{1}$ and $-\omega_{\bf b}=w_{1}\wedge w_{2}+w_{3}\wedge w_{4}\cdots w_{n-2}\wedge w_{n-1}$ and it of rank $n-1$ if $b_{1}\cdots b_{n-2}\neq 0$. Since $\dim sl_{n}^{({\bf b})}=\dim sp_{n-1}$, this implies that $sl_{n}^{({\bf b})}\cong sp_{n-1}$, and thus finishes the proof for odd $n$.
Let $n$ be even. It well-known that $\det M=\prod\limits_{i=1}^{k}m_{2i,2i-1}m_{2i-1,2i}$ for any tridiagonal $2k\times 2k$-matrix $M$ with $m_{ii}=0$.
This implies that $\omega_{\bf b}$ is of rank $n$ if $b_{1}\cdots b_{n-2}\neq 0$. Therefore, in this case, the annihilator of $\omega_{\bf b}$ in $sl_{n}$ is isomorphic to $sp_{n}$. Therefore, $sl_{n}^{({\bf b})}$ identifies with a Lie subalgebra of $sp_{n}$.
Since $V$ is a simple $sp_{n}$-module, then for any nonzero $v\in V$ the annihilator of $v$ is isomorphic to the maximal parabolic subalgebra of $sp_{n}$, i.e., to $sp_{n-1}$.
Since $sl_{n}^{(b)}$ also annihilates $v^{1}\in V$, $sl_{n}^{(b)}$ identifies with a Lie subalgebra of $sp_{n-1}$. Then the dimension argument once again implies that $sl_{n}^{({\bf b})}\cong sp_{n-1}$ as well for even $n$.
The theorem is proved. $\square$
3.5. Proof of Theorems 1.6, 2.3, and 1.10
Note that Theorem 1.6 is a particular case of Theorem 2.3(a).
Let $\hat{\hat{R}}=\mathbb{C}[a_{i},i\in I]$ an let $\hat{R}$ be its subalgebra generated by all $b_{ij}:=a_{ji}a_{i}a_{j}$ whenever $a_{ij}=-1$ in the assumptions of Theorems 1.10 and 2.3. Clearly, $\hat{R}=\mathbb{C}[b_{ij}]$.
Indeed, let $u_{i}=u_{i}^{\bf b}\in\hat{R}\otimes\mathfrak{g}$, $i\in I$ be the elements as in (the right hand sides of) Theorems 2.3 and 1.10. Let $v_{ij}^{\bf b}\in\hat{R}\otimes\mathfrak{g}$ be the left hand side of (1.3) in each case.
Lemma 3.3.
$v_{ij}^{\bf b}=0$ in each case of Theorems 1.10 and 2.3.
Proof.
Theorem 2.1 taken in $\hat{\hat{R}}\otimes sl_{n}$ and $0=k\in A_{n-1}$ (so that $A_{n-1}$ is the rooted conical tree with sub-trees $\{1,\ldots,k-1\}$ and $\{k+1,\ldots,n-1\}$) implies Lemma 3.3 in the assumptions of Theorem 2.3 for $\hat{R}\otimes sl_{n}$.
Likewise, Theorem 2.1 taken in $\hat{\hat{R}}\otimes so_{2n+1}$ and $0=n\in B_{n}$ (so that $B_{n}$ is the rooted conical tree with a single sub-tree $\{1,\ldots,k-1\}$) implies Lemma 3.3 in the assumptions of Theorem 1.10(a) in $\hat{R}\otimes so_{2n+1}$.
Finally, Theorem 2.1 taken in $\hat{\hat{R}}\otimes sp_{2n}$ and $0=n-1\in C_{n}$, $J_{+}=\{n\}$ (so that $C_{n}$ is the rooted conical tree with a single sub-tree $\{1,\ldots,n-2\}$ and $J_{+}$) implies Lemma 3.3 in the assumptions of Theorem 1.10(b) in $\hat{R}\otimes sp_{2n}$ (with $b_{i}=b_{i-1,i}$ for $i<n$ and $b_{n}=b_{n,n-1}$).
The lemma is proved. $\square$
Furthermore, we proceed similarly to the proof of Theorem 1.3. Indeed, let $R$ be any $\mathbb{C}$-algebra containing all $b_{ij}$.
Denote $A:=R\otimes U(\mathfrak{g})$. By definition, the subalgebra $B$ of $A$ generated by $X:=\{u_{i},i\in I\}$ over $R$ is $U(\mathfrak{g}^{(\bf b)})$.
The canonical specialization homomorphism $\hat{R}\twoheadrightarrow R$ implies that Lemma 3.3 holds over $R$ as well, that is, $u_{i}$ satisfy the relations of Theorems 1.10 and 2.3, that is, in the notation of Proposition 4.1, the kernel of the canonical homomorphism $\varphi_{X,B}:R\langle X\rangle\twoheadrightarrow B$ contains the set $Y:=\{v_{ij}^{\bf b},i,j\in I\}\ .$
Let us define a filtration on $A$ same way as in the proof of Theorem 1.3.
In particular, the subalgebra $B_{0}$ of $gr~{}A$ generated by $e_{i},i=1,\ldots,n-1$ is naturally isomorphic to $R\otimes U(\mathfrak{n})$, where $\mathfrak{n}=<e_{i},i\in I>$ is the nilpotent Lie subalgebra of $\mathfrak{g}$.
Then the natural inclusion of $R$-algebras $B\subset A$ is compatible with the filtration and defines a natural inclusion of graded algebras $grB\subset B_{0}$ because $gr\,u_{i}=e_{i}$ for $i\in I$ (that is, $grX=\{e_{i},i\in I\}$).
Clearly, $grY=\{(ad~{}e_{i})^{1-a_{ij}}(e_{j}):i\neq j\}$ and is the presentation of $B_{0}$, i.e., $grY$ generates the kernel of the canonical homomorphism $\varphi_{grX,grB}:R\langle grX\rangle\twoheadrightarrow grB=B_{0}$.
Thus, all conditions of Proposition 4.1 are met. Thus, the proposition guarantees that the relations of Theorems 1.10 and 2.3 give a presentation of $B=U(\mathfrak{g}^{(\bf b)})$ as an associative algebra over $R$ hence, of the Lie algebra $\mathfrak{g}^{(\bf b)}$ over $R$.
Theorems 1.6, 2.3, and 1.10 are proved. $\square$
3.6. Proof of Theorem 1.13
Proof.
The generators of $sp^{({\bf b})}_{2n}$ are as in Theorem 1.10
$u_{i}=\tilde{e}_{i}+b_{i-1}\tilde{f}_{i-1}-\delta_{i,n}\frac{1}{2}b_{n-1}^{2}[\tilde{f}_{n-1},[\tilde{f}_{n-1},\tilde{f}_{n}]]$
(a) It is clear that the set of elements
$\{u_{1},\ldots,u_{n-1}\}$
generates a subalgebra of $sp^{({\bf b})}_{2n}$ isomorphic to $sl_{n}^{(b_{1},\ldots,b_{n-2})}$. Define $J$ as an ideal of $sp^{({\bf b})}_{2n}$ generated by the set $[sl_{n}^{(b_{1},\dots,b_{n-2})},u_{n}]$.
Recall that ${sp}_{2n}$ contains $sl_{n}$ and splits as a $sl_{n}$ module
$sp_{2n}\cong V_{1}\oplus sl_{n}\oplus V_{2}$
where the $sl_{n}$ modules $V_{1}$ and $V_{2}$ are $\mathbb{C}e_{n}+[sl_{n},e_{n}]$ and $\mathbb{C}f_{n}+[sl_{n},f_{n}]$ respectively. Note that $V_{1}=S^{2}V_{\omega_{1}}$. As a direct calculation shows the projection of $J$ to $V_{1}$ is an isomorphism of vector spaces and moreover it is an $sl^{b_{1},\ldots b_{n-2}}_{n}$ module isomorphism where $V_{1}$ is a module over $sl^{b_{1},\ldots b_{n-2}}_{n}$ since $sl^{b_{1},\ldots b_{n-2}}_{n}\subset sl_{n}$.
(b) Introduce the set of elements $u^{\prime}_{i}$, $1\leq i\leq n$, as follows:
$u^{\prime}_{n}=u_{n},\,\,u^{\prime}_{n-1}=[u_{n-1},[u_{n-1},u_{n}]]$
and
$u^{\prime}_{n-k}=Ad^{2}_{u_{n-k}}\circ Ad^{2}_{u_{n-k+1}}\circ\dots\circ Ad^{2}_{u_{n-1}}(u_{n})$
There is an isomorphism of Lie algebras $\phi:sl_{n+1}^{(\bf b^{\prime})}\rightarrow J$ such that $\phi(v_{i})=u^{\prime}_{i}$, where $v_{i}$ are the generators of $sl_{n+1}^{({\bf b^{\prime}})}$ and ${\bf b^{\prime}}=(b^{\prime}_{1},\ldots,b^{\prime}_{n-1})$
$b^{\prime}_{k}=-2^{2(n-k)+1}\prod_{i=k}^{n-1}b^{2}_{i}$
$k=1,\ldots n-1$.
Moreover $[u^{\prime}_{n-k},u^{\prime}_{n-k\pm 1}]=(-2)^{k}\left(\prod_{i=1}^{k-1}b_{n-i}\right)[u_{n-k},u^{\prime}_{n-k\pm 1}]$ for $11\leq i\leq n-$ 1 and hence
$[u^{\prime}_{n-k}+(-2)^{k+1}\left(\prod_{i=1}^{k}b_{n-i}\right)u_{n-k},u^{\prime}_{n-k\pm 1}]=0$
Observe finally, that the elements
$u^{\prime\prime}_{n-k}=u^{\prime}_{n-k}+(-2)^{k+1}\left(\prod_{i=1}^{k}b_{n-i}\right)u_{n-k}$
form a copy of $sl_{n}^{(b_{1},\ldots,b_{n-2})}$ after the appropriate localization.
This finishes the proof.
$\square$
3.7. Proof of Theorems 2.7 and 2.8
We need the following fact.
Proposition 3.4.
Let $\mathfrak{g}$ be a Kac-Moody Lie algebra with $I=\{1,\ldots,r\}$ and $u_{i}:=e_{i}+b_{i-1}f_{i-1}$ (with $b_{0}=0$). Then for $k\geq 2$ one has:
(a) $(ad~{}u_{1})^{k}(u_{2})=(ad~{}e_{1})^{k}(e_{2})-2\delta_{k,2}b_{1}u_{1}$,
$(ad~{}u_{2})^{k}(u_{1})=(ad~{}e_{2})^{k}(e_{1})-2b_{1}\delta_{k,2}(u_{2}+(a_{12}+1)e_{2})$.
(b) $(ad~{}u_{i})^{k}(u_{i+1})=(ad~{}e_{i})^{k}(e_{i+1})+b_{i-1}^{k}b_{i}(ad~{}f_{i-1})^{k}(f_{i})-2b_{i}\delta_{k,2}(u_{i}-b_{i-1}(a_{i,i-1}+1)f_{i-1})$,
$(ad~{}u_{i+1})^{k}(u_{i})=(ad~{}e_{i+1})^{k}(e_{i})+b_{i}^{k}b_{i-1}(ad~{}f_{i})^{k}(f_{i-1})-2b_{i}\delta_{k,2}(u_{i+1}-b_{i+1}(a_{i,i+1}+1))$.
(c) $[u_{i},u_{j}]=0$ whenever $|i-j|>1$.
Proof.
Prove (a). Indeed,
$[u_{1},u_{2}]=[e_{1},e_{2}+b_{1}f_{1}]=e_{12}+b_{1}h_{1}$
$(ad~{}u_{1})^{2}(u_{2})=[e_{1},e_{12}+b_{1}h_{1}]=(ad~{}e_{1})^{2}(2_{2})-2b_{1}e_{1}=(ad~{}e_{1})^{2}(2_{2})-2b_{1}u_{1}$
$(ad~{}u_{2})^{2}(u_{1})=[e_{2}+b_{1}f_{1},e_{21}-b_{1}h_{1}]=(ad~{}e_{2})^{2}(e_{1})+b_{1}a_{12}e_{2}+b_{1}a_{12}e_{2}-2b_{1}^{2}f_{1}$
$=(ad~{}e_{2})^{2}(e_{1})-2b_{1}u_{2}+2b_{1}(a_{12}+1)e_{2}$
In particular,
$(ad~{}u_{1})^{3}(u_{2})=(ad~{}e_{1})^{3}(e_{2})\ ,$
$(ad~{}u_{2})^{3}(u_{1})=(ad~{}e_{2})^{3}(e_{1})+b_{1}[f_{1},(ad~{}e_{2})^{2}(e_{1})]=(ad~{}e_{2})^{3}(e_{1})\ .$
By taking commutators repeatedly with $u_{1}$ and $u_{2}$ respectively, we finish the proof of (a).
Prove (b). $[u_{i},u_{j}]=[e_{i}+t_{k}f_{k},e_{j}+t_{\ell}f_{\ell}]=e_{ij}+t_{k}t_{\ell}f_{k\ell}+\delta_{i\ell}t_{i}h_{i}-\delta_{jk}t_{j}h_{j}$, where $e_{ij}:=[e_{i},e_{j}]$, $f_{k\ell}=[f_{k},f_{\ell}]$. Then
$(ad~{}u_{i})^{2}(u_{j})=[e_{i}+t_{k}f_{k},e_{ij}+t_{k}t_{\ell}f_{k\ell}+\delta_{i\ell}t_{i}h_{i}-\delta_{jk}t_{j}h_{j}]$
$=e_{iij}+t_{k}^{2}t_{\ell}f_{kk\ell}+\delta_{i\ell}t_{i}t_{k}a_{ik}f_{k}-\delta_{i\ell}t_{i}(2e_{i}-a_{ik}t_{k}f_{k})+\delta_{jk}a_{ji}t_{j}e_{i}+\delta_{jk}t_{j}(a_{ji}e_{i}-2t_{j}f_{j})$
$=e_{iij}+t_{k}^{2}t_{\ell}f_{kk\ell}+t_{k}(2\delta_{i\ell}t_{i}a_{ik}-2\delta_{jk}t_{j})f_{k}+2(\delta_{jk}a_{ji}t_{j}-\delta_{i\ell}t_{i})e_{i}$
if $k\neq i$, $\ell\neq j$ where $e_{ab}:=[e_{a},e_{b}]$, $f_{ab}:=[f_{a},f_{b}]$, $e_{aab}:=[e_{a},e_{ab}]$, $f_{aab}:=[f_{a},f_{ab}]$.
Therefore, we obtain
$(ad~{}u_{i})^{3}(u_{j})=(ad~{}e_{i})^{3}(e_{j})+t_{k}^{3}t_{\ell}(ad~{}f_{k})^{3}(f_{\ell})$
because $[e_{i},f_{kk\ell}]=0$, $[f_{k},e_{iij}]=0$.
By taking commutators repeatedly with $u_{i}$, we obtain by induction in $r$:
$(ad~{}u_{i})^{r}(u_{j})=(ad~{}e_{i})^{r}(e_{j})+t_{k}^{r-1}t_{\ell}(ad~{}f_{k})^{r}(f_{\ell})\ .$ (3.6)
Taking $j=i+1,k=i-1,\ell=i$, $r\mapsto k$, we finish the proof of the first part of (b). Taking $i\mapsto i+1,j=k=i,\ell=i-2$, $r\mapsto k$, we finish the proof of the first part of (b) (all $t_{m}=b_{m}$).
This finishes the proof of (b).
Part (c) is immediate.
The proposition is proved. $\square$
Now we can finish proof of Theorem 2.7. Indeed, in view of Proposition 3.4(a) one can take $a^{\prime}_{12}=a_{12}$, $a^{\prime}_{21}=a_{21}$
Also, if $i>1$, then in view of Proposition 3.4(b), one can take $a^{\prime}_{i,i+1}=\min(a_{i,i+1},a_{i-1,i})$ and $a^{\prime}_{i+1,i}=\min(a_{i+1,i},a_{i,i-1})$.
Finally, in view of Proposition 3.4(c), one can take $a^{\prime}_{ij}=a_{ij}=0$ whenever $|i-j|>1$.
Thus, Theorem 2.7 is proved. $\square$
In the notation of Theorem 2.8, denote
$u^{\prime}_{i}:=\begin{cases}e_{i}+a_{i,i-1}a_{i-1}a_{i}f_{i-1}&\text{if $1\leq i\leq r$}\\ e_{i}+a_{ir}a_{r}a_{i}f_{r}&\text{if $r<i\leq r+d$}\\ \end{cases}\ .$
Clearly, the relations for the pairs $(u^{\prime}_{i},u^{\prime}_{i+1})$, $i=1,\ldots,r-1$ follow from Theorem 2.7.
It is also clear that $[u^{\prime}_{i},u^{\prime}_{j}]=0$ whenever $i=1,\ldots,r$, $j=r+1,\ldots,r+d$ and $r<i,j\leq r+d$.
Theorem 2.8 is proved. $\square$
3.8. Proof of Theorems 1.11 and 2.9
We retain the notation of Section 3.5. Theorem 2.8 implies that that the homomorphism of Theorem 2.9 is well-defined. Applying the argument of Section 3.5 once again, we obtain the desired assertion.
Thus, Theorem 2.9 is proved. $\square$
Now prove Theorem 1.11.
Taking $\mathfrak{g}=\mathfrak{g}^{\prime\prime}=\mathfrak{g}$ in Theorem 2.9 with $r=2$ we obtain the assertion of (a).
Taking $\mathfrak{g}=\mathfrak{g}^{\prime\prime}=so_{2n}$ in Theorem 2.9 with the branch at the vertex $r=n-2$, $d=2$ of the Dynkin diagram $I=\{1,\ldots,n\}$ we obtain the assertion of (b).
Prove (c). Let $\mathfrak{g}=\widehat{sl}_{n}$, $I=\{1,\ldots,n\}$, $1$ and $n$ are neighbors. Taking $sl_{n}\subset\mathfrak{g}$ corresponding to $I\setminus\{n\}=\{1,\ldots,n-1\}$ gives us relations of $sl_{n}^{({\bf b})}$ for $u_{1},\ldots,u_{n-1}$. The remaining relations of $\mathfrak{g}^{({\bf b})}$ can be obtained by the cyclic symmetry $i\mapsto i+1\mod n$.
Applying the argument of Section 3.5 once again, we see that the subalgebra generated by $u_{i}$, $i\in I$ is isomorphic to $\mathfrak{g}^{({\bf b})}$, i.e., is its edge model.
Theorem 1.11 is proved. $\square$
4. Appendix: Flat deformations
For any ring $R$ and any set $X$ denote by $R\langle X\rangle$ the free $R$-algebra freely generated by $X$.
Let $B$ be an algebra over $R$ and $X\subset B$ be a generating set. Denote by $\varphi_{X,B}$ the canonical surjection $R\langle X\rangle\twoheadrightarrow B$, denote $J_{X}:=Ker~{}\varphi_{X,B}$. The natural filtration on $R\langle X\rangle$ with $\deg x=1$ for all $x\in X$ induces a natural filtration on $B$, so let $grB$ be the associated graded algebra of $B$. Clearly, $grR\langle X\rangle=R\langle X\rangle$. For any subset $S\subset B$ denote by $grS$ the image of $S$ in $grB$.
If $A$ is a filtered algebra and $J$ is a two-sided ideal of $A$, then
$grA/J$ is canonically isomorphic to $grA/grJ$. In particular,
$grB=R\langle grX\rangle/grJ_{X}\ .$
Denote by $\varphi_{grX,grB}$ the canonical surjection $R\langle grX\rangle\twoheadrightarrow grB=R\langle grX\rangle/grJ_{X}$.
Proposition 4.1.
Let $B$ be an algebra over $R$ generated by $X\subset B$. Suppose that $Y$ is a subset of $R\langle X\rangle$ such that
$\bullet$ $Y$ is contained in the kernel of $\varphi_{X,B}$.
$\bullet$ The kernel of $\varphi_{grX,grB}$ is generated by $grY$.
Then $B=R\langle X\rangle/\langle Y\rangle$, where $\langle Y\rangle$ denotes the two-sided ideal of $R\langle X\rangle$ generated by $Y$.
Proof.
Indeed, for any subset $Y\subset R\langle X\rangle$, the inclusion $grY\subset gr\langle Y\rangle$ implies the inclusion of ideals $\langle grY\rangle\subset gr\langle Y\rangle$. Now let $Y$ be as in the assumptions of the proposition. Then the inclusion $Y\subset J_{X}$ implies the inclusion of ideals $\langle Y\rangle\subset J_{X}$ hence $gr\langle Y\rangle\subset grJ_{X}$ (if the latter inclusion is an equality, then so is the former). Taking into account that $grJ_{X}=\langle grY\rangle$, we obtain $gr\langle Y\rangle\subset\langle grY\rangle$. These two opposite inclusions imply that $gr\langle Y\rangle=grJ_{X}$. Therefore, $J_{X}=\langle Y\rangle$.
The proposition is proved. $\square$
Acknowledgments
For the third author this work is an output of a research project implemented as part of the Basic Research Program at the National Research University Higher School of Economics (HSE University). AB and VG are grateful to the MPI in Bonn, Heidelberg University, and University of Geneva where a part of this work was accomplished.